OpenStudy (anonymous):

What is the factorization of 2x2 + 7x + 6?

OpenStudy (anonymous):

2x^2+4x+3x+6 can u do it from there?

OpenStudy (anonymous):

no not at all

OpenStudy (anonymous):

LOL ok :P so basically what u have to do is the following: m*n=a*c m+n=b

OpenStudy (anonymous):

B. (2x + 2)(x + 3)

OpenStudy (anonymous):

m and n are unknown numbers, while a=2 c=6 and b=7

OpenStudy (anonymous):

um ya u got it :P

OpenStudy (anonymous):

(2x+3)(x+2)